40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10102     Entry time: Wed Jun 25 17:13:10 2014
Author: Nichin 
Type: Update 
Category: Electronics 
Subject: Laser power check - PDFR system 

[Nichin, Manasa]

I wanted to make sure Alex's system of Diode laser + laser controller + optical splitter was working fine and then make a manual measurement for AS55 PD. Manasa was supervising my work and helping me with unhooking the fibers and taking power meter readings. I have tuned on the power to REF DET from under the POY table.

I switched on the laser sitting in the 1Y1 rack and turned up the driving current to 240mA. On checking the laser power readings at AS55 (AS table) and REF DET (POY table) simultaneously, we got readings of 1.6mA and 2.4mA respectively. This much difference in readings was not expected and I did not continue taking the readings for transimpedence measurement.

I will rectify if this unequal splitting of power by the 1x16 optical splitter is going to cause any difficulties for the automated PDFR system measurement technique and resolve it if needed.

 

ELOG V3.1.3-